A motorboat heads upstream on a river that has a current of 3 miles per hour. The trip upstream takes 5 hours, and the return trip takes 2.5 hours. What is the speed of the motorboat. (the boat maintains a constant speed relative to the water)

Thanks. Hilary.

Answer by stanbon(48552)

Let speed of the boat in still water be "b".
Upstream Data:
rate = b - 3 ; time = 5 hrs.
Downstream Data:
rate = b + 3 ; time = 5/2 hrs.
EQUATION Data:
Distance up = distance down and d=rt
(b-3)5 = (b+3)(5/2)
Multiply thru by "2" to get:
10(b-3) = 5(b+3)
10b-30 = 5b+15
5b = 45
b = 9 mph (speed of boat in still water)
Cheers,
stan H